CD^{32} connection hacks

Daniel Pfund

April 10, 1995

Abstract

This text file describes:

• How to connect the CD^{32} with your Amiga;

• How to connect the CD^{32} with an Amiga 2000 style keyboard. But before you read further, please note that you still do need a CD-ROM to be able to boot the CD^{32} and communicate with another computer. In my case I used the Network CD from Almathera.

1 Disclaimer

I (Daniel Pfund) wrote this file purely to help other people to connect their CD^{32} with other peripherals. I am glad to help, but I am in no way responsible if you destroy/dammage your CD^{32} while attempting to make these connections. I have tried to keep the information as accurate as possible, but if you find an error, please contact me immediately! (see section 5 “Author” for the addresses where you can reach me).

Also please note that although this is no way a major electronics project, you should at least know how to solder correctly! Don’t attempt this if you have never soldered before …

2 Introduction

As you read in the abstract, this file will describe how to connect several items to your CD^{32} . I wanted to list here some other references you might find useful (that’s where I got the information from).

1. The CD^{32} FAQ. You can get it by FTP at: “ftp.demon.co.uk:/pub/amiga/ info/cd32-faq.txt”. This is really the main reference and you will find a huge amount of information here. The only problem being that it is very difficult to access the FTP site, even at non prime time hours.

2. Article posted by Anders Stenkvist (ask me@elixir.e.kth.se) called “CD^{32} expansion port info”. It has all the pinouts you need, but contains at least one error (the one I’ve found!), so I don’t know if the rest is correct …

3 Parts needed

The following parts are needed:

• 1 MAX 232

• 1 16 pin IC socket (optional)

• 4 1 μF elctrolytic capacities (or 10 μF goes also) rated 16V

• 1 mini DIN 6 poles male

• 1 DIN 5 poles female

• 1 DB-25 female

• 2 m (more or less) of 2 conductors (+shield) wire

• 20 cm (more or less) of 5 conductors (+shield) wire

• 1 led (optional, see section 4)

• 1 330Ω (or 470Ω) resistor (optional)

• 1 box to put the whole thing into (optional…) All the parts should be available at your local electronics store and the only costly part is the MAX 232 (∼ 6$), and all the parts would cost you about 12$.

4 The cable

Once you have got all the parts you need, you can start putting them together following the accompanying schematic in the IFF file.

The way you do this is entirely up to you. All you have to do is connect together the pins with the same name. You can do it very easily on a piece of veroboard or completely professionaly on a circuit. I’ve not provided the PCB plan, but it is not very difficult to invent one.

As you can see, you can add a led to check if the CD^{32} is on and everything is working fine. This is absolutely not mandatory. You will just need an extra led and a resistor to limit the current going through it.

5 Author

You can contact me via snail mail at:

Daniel Pfund 19, ch. de Tirelonge CH-1213 Onex Switzerland

You can also reach me via E-Mail (preferred !): “Pfund@uni2a.unige.ch” or if this address doesn’t work any more (in a couple of years) you can still write to me at: “hb9vbc@pccr01.cern.ch”. (Note the zero, not an o !)